# AutoRev: Automatic Peer Review System for Academic Research Papers

### v1

@@ Appendix B GNN Implementation Details
We trained and tested our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
B.1 Model Architecture
@@
The second GATConv layer maps the $128\times 4$ input to 256-dimensional features using 2 attention heads (output dimension $256\times 2$). A residual connection similarly projects the prior layer’s output to $256\times 2$, followed by ELU and dropout ($p=0.2$). The final GATConv layer upscales the $256\times 2$ features to the target output dimension (1024 in our case) without residual connections and dropout layers.
@@
  • A linear transformation projecting the input to a 512-dimensional space.
@@ Appendix C Finetuning Implementation Details
Table 9: Hyperparameters Used During Finetuning
Prompt for Finetuning
The prompt used for finetuning both Llama 3.1 8B and Mistral-Instruct-v0.2 can be found below:
Prompt Used During Finetuning Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction:
Generate a structured review for the research
paper passages provided below. The review should include a summary of the paper, its strengths, weaknesses, and questions for the authors. Consider that the paper is being reviewed for the ICLR conference. ### Research Paper Passages:
### Review for the paper:
@@
The extracted relevant passages are inserted ahead of the "Research Paper Passages" heading. Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
Appendix D LLM-as-a-judge evaluation
** Instruction ** : You are an expert meta-reviewer. Evaluate the following peer review text according to the scoring
[…]
Give rating for the categories described below on the full scale from 1(lowest rating) to 5(highest rating).
Assign scores independently per category (they may differ significantly).
**Category 1: Reviewer Confidence**
1 – Unqualified: Admits lack of expertise | Fundamental misunderstandings | Fails to engage with technical content
2 – Limited: Surface-level comprehension | Misses >2 key concepts | Confuses established methods
3 – Functional: Grasps main contributions | Identifies 1-2 technical flaws | Limited literature context
4 – Qualified: Precise technical critique | Contextualizes in recent work (>3 citations) |
Identifies methodological tradeoffs
5 – Authoritative: Demonstrates field leadership | Predicts follow-up research directions |
Reveals fundamental limitations in approach
**Category 2: Review Thoroughness**
1 – Negligent: <3 substantive comments | Skips methodology/results | Fails to address claims
2 – Patchy: Only addresses obvious aspects | 0 analysis of experiments | No ethical consideration
3 – Standard: Checks main claims | Basic methodology review | 1 reproducibility concern noted
4 – Rigorous: Line-by-line methodology check | Statistical validity analysis | 2+ improvement suggestions
5 – Forensic: Alternative approach proposals | Re-analysis of key results |
Compliance checklist (ethics, reproducibility, etc.)
**Category 3: Constructiveness**
1 – Destructive: Personal attacks | Unsubstantiated dismissal | 0 actionable feedback
2 – Demanding: Unrealistic requests ("redo all experiments") | Vague quality complaints
3 – Directive: General improvement areas | 1-2 specific examples | Mixed positive/negative tone
4 – Strategic: Priority-ranked suggestions | Alternative methodology paths | Conference-specific improvements
5 – Transformative: Step-by-step revision roadmap | Code/experiment snippets provided | Before-after examples
**Category 4: Helpfulness**
1 – Misleading: Factually incorrect suggestions | Contradicts paper’s goals | Harmful recommendations
2 – Abstract: "Improve writing" without examples | "Add experiments" without design
3 – Guided: Section-specific edits | 2-3 implementable suggestions | Partial error diagnosis
4 – Impactful: Directly address weaknesses | Templates/examples provided | Clear priority levels
5 – Revelatory: Exposes fundamental flaws | Provides benchmark comparisons | Supplies missing literature
**Scoring Methodology**
For EACH category:
1. Find the HIGHEST level (5-1) where ALL criteria are FULLY satisfied (Base Level)
2. Check criteria in the NEXT HIGHER LEVEL (Base Level + 1):
   - Calculate partial credit: (Satisfied Criteria in Higher Level / Total Criteria in Higher Level)
3. Final Score = Base Level + Partial Credit
4. Round to 2 decimal places (0.33 increments for 3-criteria levels, 0.5 for 2-criteria levels)
[…]
Respond ONLY with this JSON format:
  "review_scores": {{
    "Category 1": "",
    "Category 2": "",
    "Category 3": "",
    "Category 4": ""
Figure 4: Prompt used during evaluating the quality of generated reviews with LLM as a Judge
The prompt used for LLM-as-a-judge with Gemini-2.0-Flash is shown in Figure 4. The corresponding model-generated review is inserted under the appropriate subheadings within the prompt.
To demonstrate the effectiveness and sensitivity of our evaluation prompt, we include two manually crafted sample reviews: one representing a poorly written, low-quality review, and another simulating an average review. These examples are not generated by our system but are intentionally designed to test whether the LLM-as-a-judge can distinguish between different qualities of feedback. This validation step supports the reliability of the evaluation framework in scoring model outputs.
Example 1: Structured Critical Review (Moderate Quality)
Summary: The paper tackles a topic that lacks clear novelty, presents a poorly motivated approach, and delivers incomplete experiments with unconvincing results. The writing is disorganized, and key methodological and comparative details are missing, making it difficult to assess the contribution.
Strengths:
  • Addresses a relevant area of graph-based modeling (though similar to existing work).
  • Attempts to combine embeddings and graph structure in a single framework.
Weaknesses:
  • Motivation is vague and fails to establish why the problem matters.
  • No thorough comparison to closely related methods; literature review is superficial.
  • Methodology section is under-specified, lacking algorithmic details and hyperparameter settings.
  • Experimental results are incomplete: missing error bars, statistical analysis, and ablation studies.
  • Figures and tables are low resolution or poorly labeled, hindering interpretability.
Questions:
  • What is the core novelty compared to the most recent graph-embedding methods?
  • How were key hyperparameters chosen, and how sensitive are the results to those choices?
  • Can you provide ablation studies showing the contribution of each component in the model?
  • Why are error bars and statistical significance tests omitted from the reported results?
  • How does your method scale to larger datasets, and what is its computational complexity?
The evaluation scores obtained for Example-1 are as follows,
  1. 1.
    Reviewer Confidence : 3.0
  2. 2.
    Review Thoroughness : 3.0
  3. 3.
    Review Constructiveness : 3.33
  4. 4.
    Review Helpfulness : 3.0
Example 2: Informal and Unstructured Review (Low Quality)
This paper is pretty weak. The writing is sloppy and hard to follow. I’m not convinced the idea is novel. There are lots of papers out there doing similar stuff. The experiments seem incomplete, and the results aren’t very impressive. I don’t think this will be of interest to ICLR readers. You should do more experiments or maybe scrap the whole thing and start over.
Comments:
  • The motivation is unclear.
  • There’s no thorough comparison to related work.
  • The methodology description is too brief.
  • Results tables have no error bars.
  • Figures are low resolution and hard to read.
Recommendation: Reject.
The evaluation scores obtained for Example-2 are as follows,
  1. 1.
    Reviewer Confidence : 2.0
  2. 2.
    Review Thoroughness: 1.0
  3. 3.
    Review Constructiveness : 2.0
  4. 4.
    Review Helpfulness : 2.0

### v2

@@ Appendix B GNN Implementation Details
We train and test our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
B.1. Model Architecture
@@
The second GATConv layer maps the $128\times 4$ input to 256 dimensional features using 2 attention heads (output dimension $256\times 2$). A residual connection similarly projects the prior layer’s output to $256\times 2$, followed by ELU and dropout ($p=0.2$). The final GATConv layer upscales the $256\times 2$ features to the target output dimension (1024 in our case) without residual connections and dropout layers.
@@
  • A linear transformation projecting the input to a 512 dimensional space.
@@ Appendix C Finetuning Implementation Details
C.1. Prompt for Finetuning
The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing C.1.
@@ C.1. Prompt for Finetuning
The extracted relevant passages are inserted ahead of the ”Research Paper Passages” heading. Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
Appendix D Human Evaluation
Due to space constraints, the detailed evaluation guidelines are made available in an anonymized GitHub repository: Anonymous Repository. For the camera-ready version, we will release these materials in a public GitHub repository to ensure full transparency and accessibility. We will now proceed to show the snapshots of our annotation UI, developed using Streamlit  [^10] https://streamlit.io/.
D.1. Qualitative Assessment
Examples of the user interface used for the Qualitative Assessment can be seen in Figure 4 and Figure 5.
Figure 4. This interface component corresponds to the Qualitative Assessment task described in Section 5.3.1. It marks the beginning of the review presentation. Due to space constraints, only the summary and a subset of the strengths from one review are shown here.:
Figure 5. This interface component corresponds to the task described in Section 5.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.:
D.2. Factual Validity Assessment
Examples of the user interface used for Factual Validity Assessment can be seen in Figure 6a and Figure 6b.
(a) The original paper is displayed in a PDF viewer to provide factual context.
(b) Each point of the review is rated for accuracy using a dropdown menu.
Figure 6. The user interface for the Factual Validity Assessment task described in Section 5.3.2. On the left (a), the full research paper is presented to the annotator. On the right (b), the annotator rates each point of the review based on its consistency with the source document.:
Appendix E Example of a Generated Review
We present a review generated by our system for the paper: On the Role of Edge Dependency in Graph Generative Models [^11] https://openreview.net/forum?id=LCQ7YTzgRQ
Summary
The paper introduces a novel framework for evaluating generative models of graphs, focusing on the concept of “overlap” to ensure both accuracy and edge diversity. It categorizes graph generative models into three levels of complexity: edge independent, node independent, and fully dependent models. The paper derives theoretical bounds on triangle counts and other short-length cycles based on overlap and provides new generative models for each level. However, the paper’s novelty and practical contributions are questioned due to the use of existing methods and the lack of significant empirical validation. The experimental section is seen as limited and does not convincingly demonstrate the superiority of the proposed methods over existing ones.
Strengths
  • The paper introduces a novel evaluation framework that considers the importance of model-generated graph overlap to ensure both accuracy and edge diversity, which is a significant contribution to the field of graph generative models.
  • The paper presents a hierarchy of graph generative models categorized into three levels of complexity, which encapsulates a wide range of prevalent methods and provides a structured framework for evaluating graph generative models.
  • The theoretical bounds on the number of triangles and other short-length cycles are well-presented and support the findings with instances demonstrating the asymptotic optimality of these bounds.
  • The authors introduce new generative models for each of the three hierarchical levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which is a practical application of the theoretical findings.
Weaknesses
  • The novelty of the paper is limited as it primarily relies on existing methods and theoretical frameworks, and the experimental results do not convincingly demonstrate the superiority of the proposed methods over existing ones.
  • The empirical evaluation is weak and does not provide sufficient evidence to support the claims made in the paper. The experimental results are not robust and the paper does not provide enough comparisons with other models to establish the effectiveness of the proposed methods.
  • The paper has several presentation issues, including unclear definitions and notations, which could hinder the understanding of the content for readers not familiar with graph generative models.
  • The practical applications and the significance of the proposed methods are not clearly demonstrated, and the paper lacks a detailed discussion on the limitations and potential future work.
Questions
  • Can the authors clarify the definitions and notations used in the paper, particularly the ones related to the overlap and triangle density, to ensure a clearer understanding of the content?
  • How do the authors justify the novelty of their work given the reliance on existing methods and theoretical frameworks?
  • What specific contributions does the paper make to the field of graph generative models, and can these be clearly articulated?
  • Why was the empirical evaluation conducted in a limited manner, and how can the paper be improved to provide more robust and convincing experimental results?
  • Could the authors discuss the practical applications of the proposed methods and the significance of their work in more detail?
  • Are there any plans to extend the theoretical bounds to higher-order cycles, and how might this affect the overall conclusions drawn in the paper?